OpenStudy (anonymous):

The show 60 minutes recently had a share of 20, meaning that among the TV sets in use, 20% were tuned to 60 Minutes. An independent advertiser wants to check this statistic. They begin with 15 households having TV sets and survey during the time 60 Minutes is on. i. What is the probability that exactly 8 of the households are tuned to 60 Minutes. A. 0.9966 B. 0.0035 C. 0.9992 D. 0.0138 ii. What is the probability that exactly 4 of the households is tuned to 60 Minutes? A. 0.1876 B. 0.8124 C. 0.8358 D. 0.1642 iii. Find the mean, μ, of the binomial probability distribution: A. 2.4 B. 7.5 C. 3 D.

OpenStudy (dumbcow):

Use binomial distribution P(x) = nCx*p^x*(1-p)^n-x n = total num of households x = households watching 60 min p = probability household is watching 60 min based on recent ratings Part 1: n=15 x=8 p=0.2 P(8) = 15C8*(.2)^8*(.8)^7 = 0.0035 Part 2: n=15 x=4 p=0.2 P(4) = 15C4*(.2)^4*(.8)^11 = 0.1876 Part 3: mean = n*p mean = 15*0.2 = 3 Hope this helps